OpenStudy (anonymous):

what would be the linear equation for m=3 and your givin the coordinates (-2 , -2 )

OpenStudy (anonymous):

just use formula :) y-(-2) = 3* (x-(-2)) y+2=3(x+2) y=3x+4 do you see it...?

OpenStudy (anonymous):

what was the formula again? its sort of confusing.

OpenStudy (anonymous):

y-y1 = m*(x-x1) where is x1,y1 your given (-2,-2)

OpenStudy (anonymous):

are we good?

OpenStudy (anonymous):

uhhh.... i think. im only in middle school.

OpenStudy (anonymous):

not to worry :) Good Luck! Just remember (-2,-2) means x1=-2, y1=-2 right? substitute... :)